# MultiMatch: Multihead Consistency Regularization Matching for Semi-Supervised Text Classification

### v2

@@ Abstract
We introduce MultiMatch, a novel semi-supervised learning (SSL) algorithm combining the paradigms of co-training and consistency regularization with pseudo-labeling. At its core, MultiMatch features a three-fold pseudo-label weighting module designed for three key purposes: selecting and filtering pseudo-labels based on head agreement and model confidence, and weighting them according to the perceived classification difficulty. This novel module enhances and unifies three existing techniques – heads agreement from Multihead Co-training, self-adaptive thresholds from FreeMatch, and Average Pseudo-Margins from MarginMatch – resulting in a holistic approach that improves robustness and performance in SSL settings. Experimental results on benchmark datasets highlight the superior performance of MultiMatch, achieving state-of-the-art results on 9 out of 10 setups from 5 natural language processing datasets and ranking first according to the Friedman test among 19 methods. Furthermore, MultiMatch demonstrates exceptional robustness in highly imbalanced settings, outperforming the second-best approach by 3.26% – and data imbalance is a key factor for many text classification tasks. We will make our code publicly available.
MultiMatch: Multihead Consistency Regularization Matching for Semi-Supervised Text Classification
Iustin Sirbu1, Robert-Adrian Popovici1, Cornelia Caragea2, Stefan Trausan-Matu1, Traian Rebedea1,3 1National University of Science and Technology POLITEHNICA Bucharest, 2University of Illinois Chicago, 3NVIDIA Correspondence: iustin.sirbu@upb.ro
@@ 1 Introduction
In recent years, semi-supervised learning has been dominated by two main paradigms. The first, pseudo-labeling with consistency regularization, enforces prediction invariance across different augmentations of the same sample, leveraging high-confidence predictions on weakly augmented data to generate pseudo-labels for strongly augmented counterparts. The second, co-training, mitigates error accumulation by training two networks in parallel, each providing pseudo-labels for the other.
@@
To bridge this gap, we introduce MultiMatch, a novel algorithm that incorporates a three-fold pseudo-label weighting module (PLWM). This module integrates the voting mechanism from Multihead Co-training with a self-adaptive class-wise filtering strategy that leverages both historical and current predictions. Specifically, it incorporates the Average Pseudo-Margin (APM) introduced by MarginMatch to refine pseudo-label filtering based on historical predictions, while also utilizing the self-adaptive filtering mechanism of FreeMatch, which dynamically adjusts thresholds based on the model’s current predictions. Additionally, we define a taxonomy of “not useful”, “useful & difficult”, and “useful & easy” examples, and propose a weighting mechanism based on it.
We evaluate MultiMatch against 18 baselines across five text classification datasets from the unified SSL benchmark – USB Wang et al. (2022). To replicate the class imbalance challenges commonly encountered in real-world text classification tasks, we also construct highly imbalanced versions of the five datasets. Our algorithm consistently outperforms all baselines, ranking first according to both the Friedman test and the mean error across all datasets, in both balanced and imbalanced settings.
@@ 2 Related Work
These limitations have been addressed by subsequent work in various ways. For instance, FlexMatch Zhang et al. (2022) and FreeMatch Wang et al. (2023) propose to adaptively adjust class-wise thresholds based on the estimated learning status of different classes. MarginMatch Sosea and Caragea (2023) introduces the Average Pseudo-Margin (APM) score, which tracks the historical predictions for each sample and offers a more reliable confidence estimate. Other works employ the co-training Blum and Mitchell (1998) framework. Qiao et al. (2018) aim to alleviate error accumulation by using two distinct networks that are trained simultaneously, each one producing pseudo-labels for the other. JointMatch Zou and Caragea (2023) incorporates adaptive local thresholding similar to FreeMatch and introduces a weighting mechanism for disagreement examples, which are shown to be more informative. Multihead Co-training Chen et al. (2021) addresses the limitation of having to train multiple classifiers by proposing a multihead architecture where each head uses the agreement of the other ones as pseudo-label.
MultiMatch proposes to harness the strengths of both co-training and consistency regularization within a unified SSL framework. While JointMatch has taken initial steps in this direction by introducing disagreement-based weighting in the co-training framework, our earlier work – Multihead APM Sîrbu et al. (2025) – introduced a multihead architecture combined with APM-based thresholding, showing strong results in disaster tweet classification. However, MultiMatch generalizes and extends this approach by incorporating a three-fold filtering mechanism, along with a taxonomy-driven weighting scheme for pseudo-labeled examples. To our knowledge, the integration of these techniques – multihead co-training, historical confidence-based filtering, self-adaptive thresholding, and example-level weighting – into a single, robust framework is novel to MultiMatch.
@@
SSL has gained significant attention in text classification due to its ability to leverage large-scale unlabeled data alongside limited labeled examples using various approaches. For example, MixText Chen et al. (2020) mitigates overfitting on limited data by integrating Mixup Zhang et al. (2018) to interpolate labeled and unlabeled examples in a hidden space. UDA Xie et al. (2020a) uses strong data augmentation and consistency regularization to enforce prediction stability across perturbed inputs. AUM-ST Sosea and Caragea (2022) refines self-training by filtering unreliable pseudo-labels using the Area Under the Margin Pleiss et al. (2020), while SAT Chen et al. (2022) enhances FixMatch Sohn et al. (2020) by re-ranking weak and strong augmentations based on input similarity.
@@ 3.1 Background
Specifically, for an unlabeled sample $u_{b}$, FixMatch creates two augmented versions using weak and strong augmentation functions, $\alpha(\cdot)$ and $\mathcal{A(\cdot)}$ respectively. Both $\alpha(u_{b})$ and $\mathcal{A}(u_{b})$ are passed through the model to obtain predictions $q_{b}=p_{m}(y|\alpha(u_{b}))$ and $Q_{b}=p_{m}(y|\mathcal{A}(u_{b}))$. The prediction on the weakly-augmented version is converted into a pseudo-label $\hat{q_{b}}=\arg\max_{y}(q_{b})$. FixMatch uses a predefined threshold $\tau$, such that only the high-confidence pseudo-labels that pass the filtering function $\mathbb{1}_{\tau}(q_{b})=(\max(q_{b})>\tau)$ are used in the unsupervised loss computation:
@@
$\mathbb{1}_{Free}^{(t)}=\mathbb{1}(\max(q_{b})>\tau_{t}(\hat{q_{b}}))$ (2)
@@ 3.2 MultiMatch: Multihead Consistency Regularization Matching
The pipeline for training a model with MultiMatch is shown in Figure 1. Each batch of unlabeled examples is augmented using the weak and strong augmentation functions $\alpha(\cdot)$ and $\mathcal{A}(\cdot)$, respectively. The two augmented versions are forwarded through the model to make predictions. Note that, similar to Multihead Co-training, the model is multiheaded, which enables the co-training functionalities, while the shared backbone preserves the inference time of a single model. We use three heads in our presentation of the algorithm, both for simplicity and because Chen et al. (2021) identified it as the optimal choice.
As multiple predictions are obtained, pseudo-label selection is less straightforward than in classical pseudo-labeling, unless disagreement samples are discarded – a strategy we consider suboptimal. To address this, we introduce a novel pseudo-label weighting module (PLWM), which serves three key functions: selecting pseudo-labels, filtering out unreliable pseudo-labels and weighting samples according to their estimated difficulty. These functions are achieved holistically through our three-fold approach, which integrates and refines head agreements, self-adaptive thresholding, and APM scores.
A pseudocode presentation of MultiMatch can be viewed in Algorithm 1. More details about the algorithm are provided below.
@@
2:for  $h$ in $1$ to $3$ do $\triangleright$ For each classification head.
3:     $i,j=\{1,2,3\}\setminus\{h\}$triangleright The other 2 heads are used for generating pseudo-labels.
4:     $\mathcal{L}_{s}^{h}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{H}(p_{b},p_{m}^{h}(y|\alpha(x_{b})))$triangleright Compute the supervised loss for head $h$ using the labeled data.
@@
8:     $\mathcal{L}_{u,t}^{h}=\frac{1}{\mu B}\sum_{b=1}^{\mu B}W_{Multi}^{h}\cdot H(\hat{q}_{b}^{h},Q_{b}^{h})$triangleright Compute the unsupervised loss for head $h$ using the generated pseudo-labels and weights.
@@
11:$\mathcal{L}_{s}=\sum_{k\in\{1,2,3\}}\mathcal{L}_{s}^{k}$   and   $\mathcal{L}_{u}=\sum_{k\in\{1,2,3\}}\mathcal{L}_{u}^{k}$
12:$\mathcal{L}=\mathcal{L}_{s}+w_{u}\cdot\mathcal{L}_{u}$triangleright Compute the total loss.
@@ 3.3 Pseudo-Label Weighting Module (PLWM)
If the heads of the model used to generate the pseudo-label reach agreement ($\hat{q}_{b}^{i}=\hat{q}_{b}^{j}$) and both have high confidence in terms of APM score, the sample is considered useful & easy, so it is included in the loss computation with weight $1$ , using the agreed pseudo-label. If one of the heads has high confidence and the other one has low confidence, in terms of APM score, we consider that sample to be useful & difficult, as one head is not confident in the prediction. In this case, the selected pseudo-label is the one generated by the confident head. We include it in the loss computation with a larger weight $w_{d}$. If both heads have low confidence or if both heads have high confidence but disagree on the label, we consider that sample to be not useful for the current iteration and assign it weight 00. This is equivalent to filtering out the sample. To define high confidence in terms of the APM score, we introduce a novel method for computing self-adaptive APM thresholds, which is detailed in Section 3.4. Finally, we employ an additional self-adaptive filtering based on the current confidence of each head. The computation of the weights $W_{Multi}^{h}$ for the head $h$ and the unlabeled batch $u_{b}$ is described formally in Eq. 5.
@@ 3.4 Self-adaptive APM Thresholds
By starting from a subset assumed to be correctly pseudo-labeled and choosing the $f=5^{th}$ percentile, we obtain a threshold estimate that favors correctly pseudo-labeled samples, minimizing the number of incorrect pseudo-labels that pass the threshold. On the other hand, MarginMatch starts from an incorrectly classified subset and uses the $95^{th}$ percentile value, thus potentially choosing a threshold that allows for more incorrect pseudo-labels to pass, leading to error accumulation for the model. Finally, adhering to our principle of minimizing the noisy pseudo-labels, we set a lower bound of 00 for the computed threshold.
@@
$(\displaystyle\gamma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h},c}^{(t)}=max(0,\text{perc}_{f}($6)
@@
4 Experiments
For the supervised baselines and the SSL baselines implemented in USB, we report the results from the USB GitHub page  [^1] https://github.com/microsoft/Semi-supervised-learning, which provide an improvement over those originally reported by Wang et al. (2022) by using a larger batch size. For the other 2 baselines (Multihead Co-training and MarginMatch) and for MultiMatch we add the implementations to the USB framework and report the average test error of 3 runs per experiment (using the same seeds). In addition to the mean error rate across tasks, we compute the Friedman rank of each algorithm as the mean rank over all tasks $\text{rank}_{F}=\frac{1}{m}\sum_{i=1}^{m}\text{rank}_{i}$, where $m=10$ is the number of evaluation setups and $\text{rank}_{i}$ is the rank of the algorithm in the $i$-th setup.
Imbalanced settings.
Since SSL algorithms exhibit significantly different performance in imbalanced settings compared to balanced ones, and the performance gap between algorithms is larger in imbalanced setups, we aim to integrate a class-imbalance semi-supervised learning (CISSL) method to assess the potential of bridging this gap. Given its versatility – allowing seamless integration with existing SSL algorithms – we conduct additional experiments by enhancing each tested SSL algorithm with ABC Lee et al. (2021) in imbalanced settings. Further details on CISSL and ABC are provided in the Appendix A.
@@ 4 Experiments
41.05 ±1.34 42.72 ±2.57 60.71 ±2.68 53.87 ±3.03 43.26 3.10 1
Table 2: Test error rates in highly imbalanced setups. The best results without/with ABC are highlighted in blue/purple. Imbalance 100: similar distributions for labeled and unlabeled data; Imbalance -100: reversed long-tail distribution for the unlabeled set.

### v3

# Cornelia Caragea2, Stefan Trausan-Matu1, Traian Rebedea1,3

@@ Abstract
We introduce MultiMatch, a novel semi-supervised learning (SSL) algorithm combining the paradigms of co-training and consistency regularization with pseudo-labeling. At its core, MultiMatch features a pseudo-label weighting module designed for selecting and filtering pseudo-labels based on head agreement and model confidence, and weighting them according to the perceived classification difficulty. This novel module enhances and unifies three existing techniques – heads agreement from Multihead Co-training, self-adaptive thresholds from FreeMatch, and Average Pseudo-Margins from MarginMatch – resulting in a holistic approach that improves robustness and performance in SSL settings. Experimental results on benchmark datasets highlight the superior performance of MultiMatch, i.e., MultiMatch achieves state-of-the-art results on 8 out of 10 setups from 5 natural language processing datasets and ranks first according to the Friedman test among 21 methods. Furthermore, MultiMatch demonstrates exceptional robustness in highly imbalanced settings, outperforming the second-best approach by 3.26%, a critical advantage for real-world text classification tasks. Our code is available on GitHub.
MultiMatch: Multihead Consistency Regularization Matching for Semi-Supervised Text Classificationthanks: This paper has been accepted for publication in the Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing (EMNLP 2025).
Iustin Sirbu1,4, Robert-Adrian Popovici1, Cornelia Caragea2, Stefan Trausan-Matu1, Traian Rebedea1,3 1National University of Science and Technology POLITEHNICA Bucharest, 2University of Illinois Chicago,  3NVIDIA,  4Renius Technologies iustin.sirbu@upb.ro
@@ 1 Introduction
In recent years, semi-supervised learning has been dominated by two main paradigms. The first, pseudo-labeling with consistency regularization Sohn et al. (2020); Zhang et al. (2022), enforces prediction invariance across different augmentations of the same sample, leveraging high-confidence predictions on weakly augmented data to generate pseudo-labels for strongly augmented counterparts. The second, co-training Chen et al. (2021); Zou and Caragea (2023), mitigates error accumulation by training two networks in parallel, each providing pseudo-labels for the other.
@@
To bridge this gap, we introduce MultiMatch, a novel algorithm that incorporates a three-fold pseudo-label weighting module (PLWM). This module integrates the voting mechanism from Multihead Co-training Chen et al. (2021) with a self-adaptive class-wise filtering strategy that leverages both historical and current predictions. Specifically, it incorporates the Average Pseudo-Margin (APM) introduced by MarginMatch Sosea and Caragea (2023) to refine pseudo-label filtering based on historical predictions, while also utilizing the self-adaptive filtering mechanism of FreeMatch Wang et al. (2023), which dynamically adjusts thresholds based on the model’s current predictions. Additionally, we define a taxonomy of “not useful”, “useful & difficult”, and “useful & easy” examples, and propose a weighting mechanism based on it.
We evaluate MultiMatch against 20 baselines across five text classification datasets from the unified SSL benchmark – USB Wang et al. (2022). To replicate the class imbalance challenges commonly encountered in real-world text classification tasks, we also construct highly imbalanced versions of the five datasets. Our algorithm consistently outperforms all baselines, ranking first according to both the Friedman test and the mean error across all datasets, in both balanced and imbalanced settings.
@@ 2 Related Work
These limitations have been addressed by subsequent work in various ways. For instance, FlexMatch Zhang et al. (2022) and FreeMatch Wang et al. (2023) propose to adaptively adjust class-wise thresholds based on the estimated learning status of different classes. MarginMatch Sosea and Caragea (2023) introduces the Average Pseudo-Margin (APM) score, which tracks the historical predictions for each sample and offers a more reliable confidence estimate. VerifyMatch Park and Caragea (2024) uses LLMs to assign initial pseudo-labels to unlabeled data, which are then verified with an SSL model for accuracy. Other works employ the co-training Blum and Mitchell (1998) framework. Qiao et al. (2018) aim to alleviate error accumulation by using two distinct networks that are trained simultaneously, each one producing pseudo-labels for the other. Similarly, Sadat and Caragea (2024) employ two distinct networks that are trained simultaneously but instead the networks exchange sample importance weights (not pseudo-labels). JointMatch Zou and Caragea (2023) incorporates adaptive local thresholding similar to FreeMatch and introduces a weighting mechanism for disagreement examples, which are shown to be more informative. Multihead Co-training Chen et al. (2021) addresses the limitation of having to train multiple classifiers by proposing a multihead architecture where each head uses the agreement of the other ones as pseudo-label.
MultiMatch proposes to harness the strengths of both co-training and consistency regularization. While JointMatch has taken initial steps in this direction, the combination of a multihead architecture with a self-adaptive filtering mechanism based on historical predictions remains unexplored.
@@
SSL has gained significant attention in text classification due to its ability to leverage large-scale unlabeled data alongside limited labeled examples using various approaches. For example, MixText Chen et al. (2020) mitigates overfitting on limited data by integrating Mixup Zhang et al. (2018) to interpolate labeled and unlabeled examples in a hidden space. UDA Xie et al. (2020a) uses strong data augmentation and consistency regularization to enforce prediction stability across perturbed inputs. AUM-ST Sosea and Caragea (2022); Gyawali et al. (2024) refines self-training by filtering unreliable pseudo-labels using the Area Under the Margin Pleiss et al. (2020), while SAT Chen et al. (2022) enhances FixMatch Sohn et al. (2020) by re-ranking weak and strong augmentations based on input similarity. Zou et al. (2023) propose an SSL approach that uses memory banks to handle imbalanced data. Additionally, co-training-based approaches have been explored. For example, Hosseini and Caragea (2023) train two models in parallel – one leveraging unsupervised domain adaptation and the other using SSL in the target domain – allowing them to iteratively refine each other’s predictions.
@@ 3.1 Background
Specifically, for an unlabeled sample $u_{b}$, FixMatch creates two augmented versions using weak and strong augmentation functions, $\alpha(\cdot)$ and $\mathcal{A(\cdot)}$ respectively. Both $\alpha(u_{b})$ and $\mathcal{A}(u_{b})$ are passed through the model to obtain predictions $q_{b}=p_{m}(y|\alpha(u_{b}))$ and $Q_{b}=p_{m}(y|\mathcal{A}(u_{b}))$. The prediction on the weakly-augmented version is converted into a pseudo-label $\hat{q_{b}}=\arg\max_{y}(q_{b})$. FixMatch uses a predefined threshold $\tau$, such that only the high-confidence pseudo-labels that pass the filtering function $\mathbb{1}_{\tau}(q_{b})=\mathbb{1}(\max(q_{b})>\tau)$ are used in the unsupervised loss computation:
@@
$)\mathbb{1}_{Free}^{(t)}=\mathbb{1}(\max(q_{b})>\tau_{t}(\hat{q_{b}}$)) (2)
@@ 3.2 MultiMatch: Multihead Consistency Regularization Matching
The pipeline for training a model with MultiMatch is shown in Figure 1. Each batch of unlabeled examples is augmented using the weak and strong augmentation functions $\alpha(\cdot)$ and $\mathcal{A}(\cdot)$, respectively. The two augmented versions are forwarded through the model to make predictions. Note that, similar to Multihead Co-training, the model is multiheaded, which enables the co-training functionalities, while the shared backbone preserves the inference time of a single model. We use three heads in our presentation of the algorithm, both for simplicity and because Chen et al. (2021) identified this to be the optimal choice.
As multiple predictions are obtained, pseudo-label selection is less straightforward than in classical pseudo-labeling, unless disagreement samples are discarded – a strategy we consider suboptimal. To address this, we introduce a pseudo-label weighting module (PLWM), which serves three key functions: selecting pseudo-labels, filtering out unreliable pseudo-labels and weighting samples according to their estimated difficulty. These functions are achieved through our three-fold approach, which integrates and refines head agreements, self-adaptive thresholding, and APM scores.
The pseudocode of MultiMatch can be viewed in Algorithm 1. More details about the algorithm are provided in the remainder of this section.
@@
2:for $h$ in 11 to 33 do $\triangleright$ For each classification head.
3:  $i,j=\{1,2,3\}\setminus\{h\}$ $\triangleright$ The other 2 heads are used for generating pseudo-labels.
4:  $\mathcal{L}_{s}^{h}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{H}(p_{b},p_{m}^{h}(y|\alpha(x_{b})))$ $\triangleright$ Compute the supervised loss for head $h$ using the labeled data.
@@
8:  $\mathcal{L}_{u,t}^{h}=\frac{1}{\mu B}\sum_{b=1}^{\mu B}W_{Multi}^{h}\cdot H(\hat{q}_{b}^{h},Q_{b}^{h})$ $\triangleright$ Compute the unsupervised loss for head $h$ using the generated pseudo-labels and weights.
@@
[…]
@@ 3.3 Pseudo-Label Weighting Module (PLWM)
If the heads of the model used to generate the pseudo-label reach agreement ($\hat{q}_{b}^{i}=\hat{q}_{b}^{j}$) and both have high confidence in terms of APM score, the sample is considered useful & easy, so it is included in the loss computation with weight $1$ , using the agreed pseudo-label. If one of the heads has high confidence and the other one has low confidence, in terms of APM score, we consider that sample to be useful & difficult, as one head is not confident in the prediction. In this case, the selected pseudo-label is the one generated by the confident head. We include it in the loss computation with a larger weight $w_{d}$. If both heads have low confidence or if both heads have high confidence but disagree on the label, we consider that sample to be not useful for the current iteration and assign it weight 0. This is equivalent to filtering out the sample. To define high confidence in terms of the APM score, we introduce a novel method for computing self-adaptive APM thresholds, which is detailed in Section 3.4. Finally, we employ an additional self-adaptive filtering based on the current confidence of each head. The computation of the weights $W_{Multi}^{h}$ for the head $h$ and the unlabeled batch $u_{b}$ is described formally in Eq. 5.
@@ 3.4 Self-adaptive APM Thresholds
By starting from a subset assumed to be correctly pseudo-labeled and choosing the $f=5^{th}$ percentile, we obtain a threshold estimate that favors correctly pseudo-labeled samples, minimizing the number of incorrect pseudo-labels that pass the threshold. On the other hand, MarginMatch starts from an incorrectly classified subset and uses the $95^{th}$ percentile value, thus potentially choosing a threshold that allows for more incorrect pseudo-labels to pass, leading to error accumulation for the model. Finally, adhering to our principle of minimizing the noisy pseudo-labels, we set a lower bound $\gamma_{{min}}=0$ for the computed threshold.
@@
$\gamma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h},c}^{(t)}=max({\gamma_{min}},\text{percentile}_{f}($ (6)
@@
4 Experiments
For the supervised baselines and the SSL baselines implemented in USB, we report the results from the USB GitHub page [^1] https://github.com/microsoft/Semi-supervised-learning, which provide an improvement over those originally reported by Wang et al. (2022) by using a larger batch size. For the other 4 baselines (Multihead Co-training, MarginMatch, SequenceMatch, CGMatch) and for MultiMatch we add the implementations to the USB framework and report the average test error of 3 runs per experiment (using the same seeds). In addition to the mean error rate across tasks, we compute the Friedman rank of each algorithm as the mean rank over all tasks $\text{rank}_{F}=\frac{1}{m}\sum_{i=1}^{m}\text{rank}_{i}$, where $m=10$ is the number of evaluation setups and $\text{rank}_{i}$ is the rank of the algorithm in the $i$-th setup.
Imbalanced settings.
@@ 4 Experiments
41.05 ±1.34 42.72 ±2.57 60.71 ±2.68 53.87 ±3.03 43.26 3.1 1
Table 2: Test error rates in imbalanced setups. The best results without/with ABC are highlighted in blue/purple. Imbalance 100: similar distributions for labeled and unlabeled data; Imbalance -100: reversed long-tail distribution for the unlabeled set.
Since SSL algorithms exhibit significantly different performance in imbalanced settings compared to balanced ones, and the performance gap between algorithms is larger in imbalanced setups, we integrate a class-imbalance semi-supervised learning (CISSL) method, ABC Lee et al. (2021), to assess the potential of bridging this gap. ABC accounts for class imbalance by attaching an additional balanced classifier, without discarding any data. Given its seamless integration with existing SSL algorithms, we conduct additional experiments by enhancing each tested SSL algorithm with ABC in imbalanced settings. Further details on CISSL and ABC are provided in the Appendix A.